# A non-uniform metal distribution in the ram-pressure stripped M86 group of galaxies

### v1

@@ 3.2.1 Mg/Fe increases to super-Solar towards outskirts and tail
Region (arcsec) | 400–200 (1) | 200–100 (2) | 100–24 (3) | 0–24 (4) | 100–220 (5) | 132–332 (6) | 332–460 (7) | 460–628 (8) |
(Core) | (N.E. Arm) | (Plume) | (Tail 1) | (Tail 2) | ||||
XSPEC NormGas per () | ||||||||
XSPEC NormLMXB per () | – | – | – | – | – | |||
Fe | ||||||||
Mg/Fe | ||||||||
Si/Fe | ||||||||
@@
The average Mg/Fe ratio in clusters, groups and ellipticals is (Mernier et al., 2018d). The Mg/Fe ratio in the radial region of 300-500 kpc from M87, which we assume the Mg/Fe in the ambient medium of M86, is (Simionescu et al., 2015). We see that the Mg/Fe ratio in Plume (), the group outskirts (), and the Virgo ICM () are in excellent agreement.
@@ 3.2.2 Si/Fe agrees with sub-Solar Virgo ICM ratio
We observe that, the outermost radial region (Region 1) in our analysis shows slightly sub-Solar Si/Fe ratio, in agreement with the average Si/Fe ratio in hot gas pervading the Universe (Mernier et al., 2018d). For the regions closer to the core, Si/Fe ratio drops to the Virgo ICM666In the context of this paper Virgo ICM refers to the ambient medium in Virgo Cluster where M86 voyages. values, which is . This is in contrast with the Mg/Fe distribution, which shows a clear similarity for the outermost regions with the Virgo ICM, and dissimilarity within dense regions. For Si/Fe, we see that except the very outskirt, all other regions including the tail are in agreement with the Si/Fe ratio in Virgo ICM within 1 error. Additionally, we observe that the Si/Fe ratio of the dense core is slightly lower than that of Virgo ICM, although still agrees. The North-Eastern Arm shows the highest Si/Fe ratio, however due to high uncertainty it is not possible to make any physical interpretation.
@@ 4.1 The chemical composition of the M86 galaxy core
In Fig. 9 we present the abundance ratios with respect to Fe in the M86 core. The blue circles represents the RGS and purple boxes are for EPIC measurements obtained from a box region with 0.8 arcmin size overlapping the RGS dispersion direction. In this plot, we also represent the average elemental ratios with respect to Fe in the stellar population of massive ellipticals with brown stars.
We use the stellar abundance ratios to investigate if the hot gas of M86 is depleted due to ram-pressure stripping and a new generation of hot gas is released by stellar winds as thermalised stellar mass loss products are capable of dominating the hot plasma (Mathews, 1990; Mathews & Brighenti, 2003). Chemical composition of hot gas pervading cluster member galaxies and their stellar population is considered to be decoupled. The stars in such elliptical galaxies are richer in light elements (e.g. O, Ne, Mg, Si, and S), compared to heavy Fe-peak elements (e.g. Ca, Cr, Mn, Fe and, Ni) (Conroy et al., 2013). The reason of such difference is that the light elements are mostly produced in core-collapse supernovae (SNcc), while Fe-peak elements are dominantly created in Type Ia supernovae (SNIa). Elliptical (i.e., early-type) galaxies are thought to have undergone rapid star formation early in cosmic history, with a peak around (Thomas et al., 2010), which is also referred as the "downsizing". Considering SNIa events occur later in cosmic time compared to SNcc explosions, the star-forming gas in the ISM has not been enriched by SNIa products as efficiently as SNcc products at . Therefore, the rapidly formed stars in ellipticals are richer in elements, leading super-Solar abundance ratios of elements with respect to Fe (i.e., Fe > 1 in stars). In M86, because of the ram-pressure stripping, any gas inflow with Solar abundance ratios onto the galaxy is stopped (Gunn & Gott, 1972). Therefore, any agreement with the stellar abundance ratios might indicate that due to ram-pressure stripping and the consequent accretion cut-off, the hot gas of M86 is replenished by stellar mass-loss products.
However, we see that O/Fe, Mg/Fe, and S/Fe are in agreement with the average ICM abundance ratio measurements of 44 nearby systems in the CHEERS sample (Mernier et al., 2018c), represented by yellow boxes. Overall, in the M86 core halo, we observe Solar and sub-Solar abundance ratios with respect to Fe, indicating that stellar winds with super-Solar ratios do not dominate the overall composition of the hot halo.
@@
In the next section we fit our abundance pattern to various SNe explosion models and compare the results.
4.2 Constraining the SNe models based on the abundance pattern of M86 core
Metals accumulated in the hot halos of galaxies, galaxy groups and galaxy clusters provide a unique way to test nucleosynthesis models of SNe explosions. Previous studies are mostly done for cluster scale (e.g., Werner et al., 2006; de Plaa et al., 2007; Mernier et al., 2016b; Simionescu et al., 2019), which represents the integrated yields of SNe products that are able to reached the ICM. While the closed-box ICM acts like a repository for all produced and ejected metals, individual galaxies and galaxy groups may not be closed-box systems. In order to have a complete picture of enrichment mechanisms, it is important to understand the SNe contribution variations in the lower-mass systems (for a review, see, e.g., Loewenstein, 2004). The M86 group which experiences galaxy-galaxy interaction and strong ram-pressure stripping provides a unique opportunity to constrain the nucleosynthesis models to explain the abundance pattern in such a disturbed halo.
@@ 4.2 Constraining the SNe models based on the abundance pattern of M86 core
which is the SNIa fraction to the total SNe responsible for the enrichment.
We use abunfit888https://github.com/mernier/abunfit code (Mernier et al., 2016b) to fit the abundance ratios in the M86 core with different sets of SNcc and SNIa models, and IMF slopes to obtain the best /d.o.f. ratio.
SNcc synthesis occurs when a massive star collapses onto its core, where the ejected metal yields depend on the IMF, the initial metallicity, and the equatorial rotation of the star. For SNcc, we tested (i) Nomoto et al. (2013) (hereafter Nomoto when referring to the model) model, in which the lower and upper limits are and respectively. We tested the Nomoto model with different initial metallicities . The Nomoto model assumes zero initial rotational velocity for SNcc progenitors. However, any equatorial rotation induces circulations and consequent mixing of materials within the star. This mixing leads to significant differences in the ejected abundance pattern compared to non-rotating progenitors, as discussed in detail in Sect. 4.2.1. Therefore, in addition to the non-rotating Nomoto models, we also tested (ii) Limongi & Chieffi (2018) (hereafter Limongi when reffering to the model) models which includes equatorial rotational velocity for the progenitor star. We used the Limongi models with varying initial metallicities and varying rotational velocities km/s. In Limongi models the lower and upper limits are and respectively. The Limongi models also account for stellar wind contributions to the total elemental yields before the SNcc explosions occur. The stars with collapse into black hole entirely, therefore their contribution to the enrichment is limited to the stellar winds.
@@
As for the IMF, we used the Salpater IMF (Salpeter, 1955) with , and also tested intermediate and heavy IMFs with .
The discussion based on these results for SNcc progenitors is presented in the next Sect. 4.2.1, while the SNIa models are compared in Sect. 4.2.2. Finally, the overall best-fitting results of yields of SNe and AGB stars are discussed in 4.2.3.
@@
4.2.1 SNcc explosions: How do galaxy-galaxy interaction and ram-pressure stripping affect the SNcc progenitors?
Both galaxy-galaxy interaction (Larson & Tinsley, 1978; Kennicutt et al., 1987; Mihos & Hernquist, 1996; Donzelli & Pastoriza, 1997; Moreno et al., 2019) and ram-pressure stripping (Bekki & Couch, 2003; Crowl et al., 2006; Merluzzi et al., 2016; Vulcani et al., 2018) are known to enhance the star formation, before the quenching due to gas loss. In this section, based on our results listed in Table 4.2, we discuss the possible effects of such mechanisms onto SNcc progenitors.
Stars obtain spin from the contraction of their parental molecular cloud. However, observations suggest that main-sequence stars significantly lose this angular momentum (Hartmann & Stauffer, 1989; de Mink et al., 2013) via processes such as stellar winds (Langer, 1998), transport of angular momentum from the stellar core to the envelope (Ekström et al., 2008), and magnetic braking (Skumanich, 1972). M86 does not show significant far-ultraviolet emission (Gil de Paz et al., 2007), indicating very little star formation. Therefore, the stellar population of the elliptical M86 is expected to lose a significant portion of its initial angular momentum over time.
However, given that M86 experienced galaxy-galaxy interaction with NGC 4438, and extreme supersonic ram-pressure stripping, it is possible that the resulting star-forming and turbulent environment may have triggered the formation of fast-rotating SNcc progenitors for a period, which might have significantly enriched the present hot halo of M86 galaxy core. Our large elemental sample with SNcc products (-elements) from the hot halo of M86 provides a great opportunity to test the effects of galaxy-galaxy interaction and supersonic ram-pressure stripping on the SNcc progenitors in an elliptical galaxy.
@@
First, we note that the abundance yields obtained from the non-rotating models of Limongi (v0 models) and the Nomoto models differ. Limongi integrates over wider range with more massive stars (described in Sect. 4.2), and also uses different stellar evolution code compared to Nomoto, which produces higher Fe yields due to less fallback (Nomoto et al., 2013; Limongi & Chieffi, 2018). In Fig. 10, we present the Limongi and Nomoto yields comparison, using Salpeter IMF, initial metallicity , same Seitenzahl SNIa delayed-detonation model with 1 ignition spot, without equatorial rotation. We observe that the Nomoto model provides the better statistics over Limongi. The Nomoto model estimates 18.3% SNIa contribution, while Limongi results in 9.6% SNIa contribution.
@@
Second, we observe that when no equatorial rotation is introduced, both Nomoto and Limongi provide better statistics when the 3D delayed-detonation Seitenzahl models are used over 3D pure deflagration models of Fink for SNIa explosions (see Table 4.2). However, when rotation is applied with Limongi models, the Seitenzahl and Fink models result in comparable statistics in these model combinations. The differences between SNIa models of Seitenzahl and Fink are further discussed in Sect. 4.2.2.
@@
Third, when the equatorial rotation is introduced, the best-fits provide poorer statistics in Limongi models. We see that, in both Seitenzahl and Fink combinations, the non-rotating models provide the best statistics and 150 km/s models provide slightly better statistics over 300 km/s ones. This result suggest that, although the galaxy experienced galaxy-galaxy interactions and supersonic ram-pressure stripping, these turbulent and potentially star-forming environments might not result in rotating SNcc progenitors.
@@
Last but not least, our results suggest that the best fits to the Limongi models are obtained when assuming zero initial metallicity () across all model combinations. Consistent with a zero-metallicity environment—and the resulting inefficient cooling and limited fragmentation—we observe a heavy () IMF for all SNcc progenitor stars. This result from Limongi is in tension with our findings based on the Nomoto models, which provide better statistical fits over Limongi, and instead suggest either low () or Solar () initial metallicities, along with either intermediate-heavy () or Salpeter () IMFs. In any case, our results support neither rotating SNcc progenitors nor high-metallicity initial environments. The hot gas in M86 may have been significantly enriched by an early generation of SNcc explosions originating from a low-metallicity environment, as expected for a massive elliptical galaxy that experienced rapid star formation early in cosmic history at . We find that the effects of galaxy-galaxy collisions and ram-pressure stripping on SNcc progenitors are limited.

### v2

@@ 3.2.1 Mg/Fe increases to super-Solar towards outskirts and tail
| Region (arcsec) | 400–200 (1) | 200–100 (2) | 100–24 (3) | 0–24 (4) | 100–220 (5) | 132–332 (6) | 332–460 (7) | 460–628 (8) | ||
| (Core) | (N.E. Arm) | (Plume) | (Tail 1) | (Tail 2) | ||||||
| Fe | ||||||||||
| Mg/Fe | ||||||||||
| Si/Fe | ||||||||||
@@
The average Mg/Fe ratio in clusters, groups and ellipticals is (Mernier et al., 2018b). The Mg/Fe ratio in the radial region of 300-500 kpc from M87, which we assume the Mg/Fe in the ambient medium of M86, is (Simionescu et al., 2015). We see that the Mg/Fe ratio in Plume (), the group outskirts (), and the Virgo ICM () are in excellent agreement.
@@ 3.2.2 Si/Fe agrees with sub-Solar Virgo ICM ratio
We observe that, the outermost radial region (Region 1) in our analysis shows slightly sub-Solar Si/Fe ratio, in agreement with the average Si/Fe ratio in hot gas pervading the Universe (Mernier et al., 2018b). For the regions closer to the core, Si/Fe ratio drops to the Virgo ICM666In the context of this paper Virgo ICM refers to the ambient medium in Virgo Cluster where M86 voyages. values, which is . This is in contrast with the Mg/Fe distribution, which shows a clear similarity for the outermost regions with the Virgo ICM, and dissimilarity within dense regions. For Si/Fe, we see that except the very outskirt, all other regions including the tail are in agreement with the Si/Fe ratio in Virgo ICM within 1 error. Additionally, we observe that the Si/Fe ratio of the dense core is slightly lower than that of Virgo ICM, although still agrees. The North-Eastern Arm shows the highest Si/Fe ratio, however due to high uncertainty it is not possible to make any physical interpretation.
@@ 4.1 The chemical composition of the M86 galaxy core
In Fig. 9 we present the abundance ratios with respect to Fe in the M86 core. The blue circles represents the RGS and purple boxes are for EPIC measurements obtained from a box region with 0.8 arcmin size overlapping the RGS dispersion direction. In this plot, we also compare our results with (i) the multi-temperature Suzaku measurements of M86 galaxy core by Hishi et al. (2017) (gray triangles), (ii) the average elemental ratios with respect to Fe in the 44 nearby systems of the CHEERS sample by Mernier et al. (2018b) (yellow boxes), (iii) the average X/Fe ratios in the stellar content of early-type galaxies by Conroy et al. (2013) (black stars). Finally, we also show the uncertainties in the Solar measurements (Lodders et al., 2009) in gray boxes.
We compare our hot gas abundance ratios with average stellar abundance ratios to investigate whether the original hot gas in M86 has been depleted by ram-pressure stripping and replaced by a new generation of hot gas produced by stellar winds, as thermalised stellar mass loss is a continuous internal source of hot gas (Mathews, 1990; Mathews & Brighenti, 2003). Chemical composition of hot gas pervading cluster member galaxies and their stellar population is considered to be decoupled. The stars in such elliptical galaxies are richer in light elements (e.g. O, Ne, Mg, Si, and S), compared to heavy Fe-peak elements (e.g. Ca, Cr, Mn, Fe and, Ni) (Conroy et al., 2013). The reason of such difference is that the light elements are mostly produced in core-collapse supernovae (SNcc), while Fe-peak elements are dominantly created in Type Ia supernovae (SNIa). Elliptical (i.e., early-type) galaxies are thought to have undergone rapid star formation early in cosmic history, with a peak around (Thomas et al., 2010), which is also referred as the "downsizing". Considering SNIa events occur later in cosmic time compared to SNcc explosions, the star-forming gas in the ISM has not been enriched by SNIa products as efficiently as SNcc products at . Therefore, the rapidly formed stars in ellipticals are richer in elements, leading super-Solar abundance ratios of elements with respect to Fe (i.e., Fe > 1 in stars). In M86, because of the ram-pressure stripping, any gas inflow with Solar abundance ratios onto the galaxy is stopped (Gunn & Gott, 1972). Therefore, any agreement with the stellar abundance ratios might indicate that due to ram-pressure stripping and the consequent accretion cut-off, the hot gas of M86 is replenished by stellar mass-loss products.
However, we see that O/Fe, Mg/Fe, and S/Fe are in agreement with the average ICM abundance ratio measurements of 44 nearby systems in the CHEERS sample (Mernier et al., 2018b), represented by yellow boxes. Overall, in the M86 core halo, we observe Solar and sub-Solar abundance ratios with respect to Fe, indicating that stellar winds with super-Solar ratios do not dominate the overall composition of the hot halo.
@@
Moreover, we observe that our results of XMM-Newton/EPIC and RGS are in agreement with the Suzaku measurements by Hishi et al. (2017), except Ne/Fe, which they found an unusually high value, greater than 3 Solar. In addition to instrumental differences, the main distinctions between our study and Hishi et al. (2017) lie in the choice of temperature models and background treatments. We adopted the vgadem model, while they used 2T models. In our analysis, we found that both approaches –vgadem and 2T– yield consistent results. For the background, we determined the parameters from a full FOV fit and rescaled the normalizations for each region to improve stability, whereas they fixed the normalizations of all background components—except the ICM—without such rescaling. Other differences include the assumed foreground temperatures and the inclusion of an SWCX component in our model, which was not included in theirs.
4.2 The dissimilar metal distribution in the hot gas of M86
In their Suzaku analysis, Hishi et al. (2017) reported Solar abundance ratios everywhere within the uncertainties, except for Ne/Fe, which they found to be super-Solar throughout. The only deviation of their results from the overall Solar ratios—aside from Ne/Fe—is their Mg/Fe ratio in the plume, for which they found a super-Solar value, consistent with ours. Their Si/Fe and S/Fe ratios for the plume are Solar, comparable with our results. Additionally, they measured a super-Solar Mg/Fe ratio in their outer group gas region at a distance from the core comparable to our Region 1, though still consistent with the Solar value within . However, for the tail, they reported Solar abundance ratios, which is inconsistent with our super-Solar Mg/Fe result, as well as our significantly sub-Solar Si/Fe and S/Fe ratios in the same tail regions. Nevertheless, our XMM-Newton results and the Suzaku results of Hishi et al. (2017) agree for the plume and core, while the tail results do not.
4.2.1 Mg/Fe ratio
The Mg/Fe ratio in Plume is in agreement with that of M86 group outskirts and Virgo ICM. The Plume structure might have formed from this aforementioned mixed low-entropy group gas after the collision with NGC 4438 and the onset of ram-pressure stripping. Since these regions extend well beyond the stellar population of the galaxy and experience an accretion cut-off from their surroundings due to ram-pressure stripping (Gunn & Gott, 1972), the chemical composition in emission structures might have remained unchanged since their formation. Similarly, the chemical composition of the extended group gas (Regions 1 and 2), which is mixed efficiently with the Virgo ICM, is expected to preserve its chemical composition compared to that of the galaxy core. In contrast, in the M86 galaxy core, SNIa events continue to enrich the medium. Therefore, SNIa products such as Fe are expected to increase in the galaxy core, leading to a decreased Mg/Fe ratio.
Additionally, ram-pressure stripping enhances the star formation, which increases the number of massive stars and SNcc explosions. In Tail 2 (Region 8), the far end of the M86 tail in our observation, we observe the highest Mg/Fe ratio in M86. It is plausible that the dynamical disruption in the M86 tail enhanced the Mg enrichment.
4.2.2 Si/Fe and S/Fe ratios
4.3 Constraining the SNe enrichment based on the abundance pattern of M86 core
Metals accumulated in the hot halos of galaxies, galaxy groups and galaxy clusters provide a unique way to test SNe enrichment scenarios. Previous studies are mostly done for cluster scale (e.g., Werner et al., 2006; de Plaa et al., 2007; Mernier et al., 2016b; Simionescu et al., 2019; Gatuzz et al., 2023b), which represents the integrated yields of SNe products that are able to reach the ICM. While the closed-box ICM acts like a repository for all produced and ejected metals, individual galaxies and galaxy groups may not be closed-box systems. In order to have a complete picture of enrichment mechanisms, it is important to understand the SNe contribution variations in the lower-mass systems (for a review, see, e.g., Loewenstein, 2004). The M86 group which experiences galaxy-galaxy interaction and strong ram-pressure stripping provides a unique opportunity to constrain the SNe enrichment in such a disturbed halo.
@@ 4.3 Constraining the SNe enrichment based on the abundance pattern of M86 core
which is the SNIa fraction to the total SNe responsible for the enrichment. Using the average abundance pattern in 44 systems from the CHEERS sample, which are mostly galaxy clusters, Mernier et al. (2016b) calculated that the SNIa contribution to the total enrichment of hot halos is approximately 29–42%. Similarly, Gatuzz et al. (2023a) recently showed that the SNIa contribution to the enrichment of the Virgo ICM, through which M86 travels, is 31–47%.
In order to find the SNIa and SNcc contribution to the total metal enrichment, we use abunfit888https://github.com/mernier/abunfit code (Mernier et al., 2016b) to fit the abundance ratios in the M86 core with different sets of SNcc and SNIa models, and IMF slopes to obtain the best /d.o.f. ratio. We test a wide range of possibilities for SNe progenitors to comprehensively investigate the SNe enrichment in the hot halo of M86.
SNcc synthesis occurs when a massive star collapses onto its core, where the ejected metal yields depend on the IMF, the initial metallicity, and the equatorial rotation of the star. For SNcc, we tested (i) Nomoto et al. (2013) (hereafter Nomoto when referring to the model) model, in which the lower and upper limits are and respectively. We tested the Nomoto model with different initial metallicities . The Nomoto model assumes zero initial rotational velocity for SNcc progenitors. However, any equatorial rotation induces circulations and consequent mixing of materials within the star. This mixing leads to significant differences in the ejected abundance pattern compared to non-rotating progenitors, as discussed in detail in Sect. 4.3.1. Therefore, in addition to the non-rotating Nomoto models, we also tested (ii) Limongi & Chieffi (2018) (hereafter Limongi when reffering to the model) models which includes equatorial rotational velocity for the progenitor star. We used the Limongi models with varying initial metallicities and varying rotational velocities km/s. In Limongi models the lower and upper limits are and respectively. The Limongi models also account for stellar wind contributions to the total elemental yields before the SNcc explosions occur. The stars with collapse into black hole entirely, therefore their contribution to the enrichment is limited to the stellar winds.
@@
As for the IMF, we used the Salpeter IMF (Salpeter, 1955) with , and also tested intermediate and heavy IMFs with .
We emphasize that, our analysis holds supernova parameters of IMF slope, progenitor metallicity, explosion geometry, and stellar rotation are fixed for all progenitors, even though in reality each of these properties varies for each star. To definitively identify explosion scenarios in M86 is out of scope of this paper. Our goal in this section is to compare how these SN models with fixed parameters that are previously constrained by abundance patterns in closed-box galaxy clusters perform when applied to an open system like M86.
@@
The discussion based on these results for SNcc progenitors is presented in the next Sect. 4.3.1, while the SNIa models are compared in Sect. 4.3.2. Finally, the overall best-fitting results of yields of SNe and AGB stars are discussed in 4.3.3.
4.3.1 SNcc explosions: How do galaxy-galaxy interaction and ram-pressure stripping affect the SNcc enrichment?
Both galaxy-galaxy interaction (Larson & Tinsley, 1978; Kennicutt et al., 1987; Mihos & Hernquist, 1996; Donzelli & Pastoriza, 1997; Moreno et al., 2019) and ram-pressure stripping (Bekki & Couch, 2003; Crowl et al., 2006; Merluzzi et al., 2016; Vulcani et al., 2018) are known to enhance the star formation, before the quenching due to gas loss. In this section, based on our results listed in Table 4.3, we discuss the possible effects of such mechanisms onto SNcc enrichment.
M86 does not show significant far-ultraviolet emission (Gil de Paz et al., 2007), indicating very little star formation. However, given that M86 experienced galaxy-galaxy interaction with NGC 4438, and extreme supersonic ram-pressure stripping, it is possible that the resulting star-forming and turbulent environment may have triggered the formation of fast-rotating SNcc progenitors for a period, which might have significantly enriched the present hot halo of M86 galaxy core.
@@
First, we note that the abundance yields obtained from the non-rotating models of Limongi (v0 models) and the Nomoto models differ. Limongi integrates over a wider range with more massive stars (described in Sect. 4.3), and also uses different stellar evolution code compared to Nomoto, which produces higher Fe yields due to less fallback (Nomoto et al., 2013; Limongi & Chieffi, 2018).
@@
Second, we observe that when no equatorial rotation is introduced, both Nomoto and Limongi provide better statistics when the 3D delayed-detonation Seitenzahl models are used over 3D pure deflagration models of Fink for SNIa explosions (see Table 4.3). However, when rotation is applied with Limongi models, the Seitenzahl and Fink models result in comparable statistics in these model combinations. The differences between SNIa models of Seitenzahl and Fink are further discussed in Sect. 4.3.2.
@@
Third, when the equatorial rotation is introduced, the best-fits provide poorer statistics in Limongi models. We see that, in both Seitenzahl and Fink combinations, the non-rotating models provide the best statistics and 150 km/s models provide slightly better statistics over 300 km/s ones.
@@
Last but not least, our results suggest that the best fits to the Limongi models are obtained when assuming zero initial metallicity () across all model combinations. Consistent with a zero-metallicity environment—and the resulting inefficient cooling and limited fragmentation—we observe a heavy () IMF for all SNcc progenitor stars. This result from Limongi is in tension with our findings based on the Nomoto models, which provide better statistical fits over Limongi, and instead suggest either low () or Solar () initial metallicities, along with either intermediate-heavy () or Salpeter () IMFs.
We caution the reader that our time-invariant IMF assumption is a simplification intended to compare trends between closed-box galaxy clusters and the relatively open M86 system. Similarly, the rotation and metallicity of SNcc progenitors are assumed to be uniform both spatially and temporally. In reality, however, the picture is more complex, as these parameters vary between stars and between different evolutionary episodes of M86. Our aim is not to determine the exact explosion scenario but rather to test a wide range of possibilities. By doing so, we show that SNcc enrichment in M86 is consistently high. We find that given the delayed-detonation SNIa explosions, which are preferred by the supernovae community, our SNIa contribution result for M86 from the best fits lies in the range of 8–25%. This contribution ratio is significantly lower than the 31–47% contribution in the Virgo Cluster (Gatuzz et al., 2023a), 29–42% of the averaged abundance sample from the CHEERS sample, consisting mostly of clusters (Mernier et al., 2016b). The higher contribution of SNcc enrichment in the hot gas of M86 suggests an increased starburst and subsequent SNcc explosions in the galaxy. Our results indicate that the ram-pressure stripping and galaxy-galaxy interaction induced a starburst episode, which led to the increased SNcc enrichment and high SNcc contribution.